OpenStudy (anonymous):

Consider the function f(x)=x+2cos(x), 0

OpenStudy (anonymous):

i would start by getting the first derivative and setting it to zero

OpenStudy (anonymous):

inorder to get the critical points

OpenStudy (anonymous):

I have both derivatives f'(x): 1-2sinx and f''(x): -2cosx Then the rest I don't know how to find.

OpenStudy (anonymous):

set them to zero and solve for x

OpenStudy (anonymous):

keeping in mind, the interval your are working within

OpenStudy (anonymous):

For the first one I got sinx=1/2 and the second 0. Is that right?

OpenStudy (anonymous):

oaky, so at what angle(s) within 0<x<2pi is sin(x), 1/2? at what angle(s) withing that interval is cos(x), 0?

OpenStudy (anonymous):

60 and 30. I don't think I'm doing this right.

OpenStudy (anonymous):

Trig throws me off.

OpenStudy (anonymous):

actually, the critical points come from the first derivative, so if you said the first derivative is 1-2sin(x)=0 sin(x)=1/2 so the occurs at 30 deg or pi/6